Circular No. 7482

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000dh IN IC 5374 T. Puckett, Mountain Town, GA; and D. George, Ottawa, ON, report

the discovery of an apparent supernova (mag 19.4) on an unfiltered CCD frame (limiting mag 20.5) taken with the Puckett Observatory 0.60-m automated supernova patrol telescope on Aug. 24.22 UT. SN 2000dh is located at α = 0^{h}01^{m}03^{s}.94, δ = +4^{o}30′05′′.4 (equinox 2000.0), which is 8′′.7 west and 4′′.7 north of the center of IC 5374. The new object was also present on an unfiltered CCD frame taken on Aug. 25.29, but it does not appear on Palomar Sky Survey images taken on 1991 Sept. 13 (limiting mag ∼ 21.0), 1992 Aug. 27 (limiting mag ∼ 21.0), or 1955 Nov. 11 (limiting mag ∼ 20.0).

EXO 1745–248 C. B. Markwardt, University of Maryland and Goddard Space Flight

Center (GSFC), NASA; and T. E. Strohmayer, J. H. Swank, and W. Zhang, GSFC, write: “The transient EXO 1745–248 in the globular cluster Terzan 5 (cf. IAUC 7454) has recently increased in intensity. RXTE PCA pointed and scanning observations from July 13–Aug. 15 show that the persistent intensity has varied erratically, increasing to 600 mCrab (2–10 keV) on Aug. 21.8 UT, which is near the Eddington limit for a source at 7600 pc. Dipping activity is also present, but no eclipse events have been seen. Fifteen x-ray bursts were detected, with peak fluxes between 200 and 350 mCrab and an average separation of 25 min. The x-ray spectrum of the bursts at the peak was consistent with blackbody emission with kT = 2.4–2.8 keV and radii of 2–4 km. Only modest cooling is suggested in the tails of the bursts. The short burst recurrence time, the range of peak luminosities, and the weak cooling suggest that the bursts may be of type II (accretion instability), but we cannot exclude a thermonuclear origin. The persistent emission of Aug. 13 is best fitted by a blackbody and power law with exponential cutoff plus an iron emission feature near 6.6 keV, typical of bright, low-mass x-ray binaries. Two quasiperiodic features were present on Aug. 13, with centroids of 65 and 134 mHz, the fundamental having a fractional r.m.s. amplitude of 10 percent. Optical observations are encouraged.”

CI AQUILAE Visual magnitude estimates by M. Lehký, Hradec Králové, Czech Re-

public: June 7.875 UT, 11.0; 10.968, 11.7; July 25.958, 12.9; 30.954, 12.8; Aug. 1.916, 12.8; 19.909, 12.9; 20.899, 12.9.

2000 August 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green